OpenStudy (anonymous):

Find inverse of y=2e^x-2

OpenStudy (anonymous):

here y+2=2*e^x (y+2)/2 =e^x taking log to the base e both sides we have x= ln((y+2)/2) when it is ln it is always understood that the base of loagrithm is e